OpenStudy (anonymous):

how do you solve -9x+3y=18

OpenStudy (marigirl):

Which variable are you making the subject of the equation? x or y?

OpenStudy (anonymous):

y=6+3x x= (18-3y)/-9

OpenStudy (anonymous):